# Discovering Hidden Visual Concepts Beyond Linguistic Input in Infant Learning

### v1

@@ Abstract
Infants develop complex visual understanding rapidly, even preceding of the acquisition of linguistic inputs. As computer vision seeks to replicate the human vision system, understanding infant visual development may offer valuable insights. In this paper, we present an interdisciplinary study exploring this question: can a computational model that imitates the infant learning process develop broader visual concepts that extend beyond the vocabulary it has heard, similar to how infants naturally learn? To investigate this, we analyze a recently published model in Science by Vong et al., which is trained on longitudinal, egocentric images of a single child paired with transcribed parental speech. We introduce a training-free framework that can discover visual concept neurons hidden in the model’s internal representations. Our findings show that these neurons can classify objects outside its original vocabulary. Furthermore, we compare the visual representations in infant-like models with those in modern computer vision models, such as CLIP or ImageNet pre-trained model, highlighting key similarities and differences. Ultimately, our work bridges cognitive science and computer vision by analyzing the internal representations of a computational model trained on an infant’s visual and linguistic inputs.
@@
Figure 1: Discover visual concepts hidden in internal representation of infant model: Infants learn visual representations not only from parental guidance but through self-exploration. Based on this, we hypothesize that a model trained on infant egocentric images and its parent-spoken words develops internal visual representations of objects that extend beyond parental speech vocabulary. (a) Prior work trains a model to imitate the infant’s word learning. (b) Our neuron labeling approach finds neurons sensitive to words that are outside the vocabulary of parental speech, thus discovering visual concept neurons (e.g.rug”). (c) Initially, the model can only classify objects within its training vocabulary. (d) These concept neurons enables the model to recognize out-of-vocabulary concepts without any model parameter updates, aligning with cognitive studies showing infants develop advanced visual concepts.
@@ 1 Introduction
Infants are remarkable learners, sparking interest across various academic disciplines to uncover their learning mechanisms. Computer vision is no exception, with researchers studying infant visual learning from various perspectives [1, 31, 33, 30]. A recent milestone is Vong et al. [38], who introduces a model trained on longitudinal egocentric videos of a single infant, captured from 6 to 25 months. This model, termed the Child’s View for Contrastive Learning (CVCL), is trained to associate transcribed parental speech with corresponding video frames in a similar manner to CLIP [32]. Learning from this unique infant dataset, CVCL has demonstrated notable object recognition abilities.
These results are consistent with findings from developmental psychology, which directly studies infants through experimental and observational research. Infants learn their first object names through direct association, connecting words they hear with objects they see. They can associate words spoken by parents with their visual referents as early as 6 months of age [3, 16]. Their early object recognition develops through a prolonged, multi-stage learning process, which requires infants significant visual experience with specific objects in the complex real world. As infants accumulate visual experience with certain objects, they become better at recognizing and categorizing these objects in varied contexts. According to a headcam study done on 8.5 to 10.5 month-old infants during feeding [8], infants’ first nouns were those objects that were most frequent in their visual fields. While infants may also be hearing the names of frequently seen objects, their early visual familiarity with these common objects might come first and even help support the process of learning their names. For example, infant has shown innate ability to recognize patterns [12]. Moreover, their development of visual concepts occurs before the emergence of verbal thought [26].
Based on the infant studies shown above, we hypothesize that a model trained on infant-like data may similarly develop visual concepts that are not explicitly represented in its linguistic training data. Specifically, while the original study [38] performs classification with the vocabulary that limited to what the infant have heard from parents (in-vocabulary), the model could have developed visual concepts extending beyond the explicit linguistic input present during training phase (out-of-vocabulary).
@@
To investigate this, we analyze the CVCL’s internal representation, using network dissection [2, 29], or more intuitively, “neuron labeling”. We then introduce NeuronClassifier, a neuron-based, training-free classification framework that leverages visual concepts identified from model’s internal representations. Our approach not only achieves higher recognition accuracy than originally reported by CVCL, but also discovers meaningful out-of-vocabulary concepts embedded within the model, which supports our hypothesis. More interestingly, the out-of-vocabulary words found within the model tend to have a higher age of acquisition for infants (Figure 6), representing higher-level concepts. This indicates that the model has developed broader visual concepts ahead of linguistic understanding, aligning with cognitive studies showing that infants develop visual concepts as preverbal thought [26].
@@
Nontheless, CVCL’s recognition ability is still limited compared to models like CLIP or ImageNet-trained models, which are widely used in computer vision as generic visual representations. This raises the question: what distinguishes these higher-performing models from CVCL? To investigate this, we analyze the internal representations of CVCL in comparison with CLIP and ImageNet models. We find that while CVCL exhibits similar low-level features in its early layers, its high-level features in the final layer differ significantly. These differences extend to the visual concept neurons present in the model’s deeper layers.
@@
  • We show that infant model have developed understanding beyond linguistic training inputs, by discovering visual concepts hidden in the model representations, aligning with cognitive naturalness.
  • We propose a training-free classification framework that leverages the model’s internal representations through neuron labeling, which helps us understand the visual representations within the infant model.
  • We find that the infant model shares similar low-level representations with large-scale pre-trained models but diverges in deeper layers due to a lack of diverse high-level visual concepts.
@@ 2 Related Work
Modeling how children learn has long been a strategy for advancing artificial intelligence. Instead of directly replicating adult intelligence, Alan Turing suggested, “why not rather try to produce one which simulates the child’s?” [37] With this in mind, training models on infant’s egocentric videos [1, 30, 33, 31, 36, 38] is a natural direction, because they are an approximation of real training data for infant vision system. This work builds upon CVCL model [38], which we explain in Sec. 2.1.
@@
Our goal is to understand the model trained on infant data. In this context, techniques for interpreting intermediate representations in deep neural networks are relevant. Beginning with Network Dissection [2], a method that quantifies alignment between hidden neurons and visual concepts, numerous studies have aimed to make black-box models more transparent. These methods enable compositional concept discovery [27], assignment of compositional concepts with statistical quantification [5], open-vocabulary neuron captioning [20], and the use of CLIP’s rich embeddings for neuron-concept alignment [29, 21]. Beyond direct neuron dissection, other approaches analyze component functions by decomposing image representations to reveal the role of attention heads within multimodal embedding spaces [14] and identifying neurons with similar functions across a diverse model zoo [11].
2.1 Preliminary: CVCL model
@@
The CVCL model is trained on the SAYCam-S dataset [35], a longitudinal collection of egocentric recordings from a child aged 6 to 25 months, containing around 200 hours of video. To create meaningful image-text pairs for model training, transcripts were pre-processed to retain only child-directed utterances, excluding the child’s own vocalizations. Frames were extracted to align with utterance timestamps. The resulting dataset comprises 600,285 frames paired with 37,500 transcribed utterances, forming a multimodal dataset that simulates the sparse and noisy real-world experiences from which children learn.
@@
Employing a self-supervised contrastive learning approach akin to CLIP [32], the CVCL model learns to align egocentric visual frames with transcribed parental speech. Co-occurring pairs are treated as positive examples, while non-co-occurring pairs serve as negatives. This method allows the model to develop multimodal representations without external labels, imitating a child’s natural learning process.
@@
For evaluation, CVCL adopts a $n$-way classification task in which the model selects the most relevant visual reference from a set comprising one target image and $n-1$ foil images. This approach is inspired by the intermodal preferential looking paradigm (IPLP) [15, 16], used in infant recognition studies to measure language comprehension through differential visual fixation. By aligning its visual and text encoders, CVCL achieves comparable in-domain test accuracy to models like CLIP. However, for out-of-domain generalization, CVCL demonstrates relatively weak performance on the Konkle object dataset [22], which includes naturalistic object categories on a white background, using only classes available in the training data.
@@ 3 Method
In this section, we describe how to explore neuron-level concepts and leveraging them in $n$ way classification tasks. We begin by using published neuron labeling techniques to discover visual concepts hidden within the CVCL model [38]. Then, we introduce our framework to utilize these labeled neuron concepts for $n$-way classification.
@@ 3.1 Neuron Labeling
We follow CLIP-Dissect [29] for internal representation analysis due to its flexibility in concept sets and probing images, which helps us easily identify visual concept neurons.
@@
$l_{k}=\arg\max_{m}\texttt{sim}(t_{m},q_{k};P).$ (3)
In here $\texttt{sim}(\cdot,\cdot)$ represents the similarity function (e.g., cosine similarity, Soft-WPMI[39, 29]). Collecting the label $l_{k}$ for each neuron, we define the label vector $\mathcal{L}=[l_{1},l_{2},\dots,l_{K}]$ to represent the assigned concepts across the entire model.
During this neuron labeling process, we aim to assign meaningful concepts to each neuron. The two main reasons for choosing CLIP-Dissect are:
  • Concept set $\mathcal{S}$: Instead of allowing an infinite range of possible concepts for neuron labeling, using a fixed concept set narrows this process by constraining it to a limited selection of concepts. Including high-level concepts in this set enables us to identify neurons corresponding to these visual concepts. This setup ensures that each neuron is assigned a specific label, though some labels may be spurious, as illustrated in Figure 3.
@@
Figure 3: Examples of Spurious Labeled Neurons obtained using CLIP-Dissect across all 4 layers in the CVCL-ResNeXt50 vision encoder. For each selected neuron, the top-3 activated images from the Konkle object dataset are displayed.
@@
The flexibility of the concept set $\mathcal{S}$ provides a spectrum of possibilities, allowing us to discover visual concepts that the infant model has never encountered in its training data. In this process, CLIP serves as a well-pretrained miner, utilizing its rich image-text embeddings to identify concepts hidden within the CVCL model. By aligning the activations of CVCL’s neurons with CLIP’s embeddings, we uncover meaningful hidden visual concepts within the infant model, even extending beyond the model’s linguistic training data. This approach leverages the diverse vocabulary of the concept set and the rich embeddings of CLIP to reveal visual concepts embedded in the CVCL model’s internal representations, each associated with corresponding neurons.
@@ 3.2 Neuron-Based Classification
How do we ensure that neurons representing concepts beyond the model’s original vocabulary truly exist within the network? In this section, we propose NeuronClassifier, a training-free framework that leverages neuron activations to detect and validate such concepts. By identifying neurons with specific visual concept, we aim to confirm the presence of these latent, beyond-vocabulary neurons and use them to perform $n$-way classification. The framework, illustrated in Figure 4, involves three main steps.
@@
Given a target label $l_{k}\in\mathcal{L}$, where $\mathcal{L}$ represents all neurons’ label in the model, same as label vector in Section 3.1 (e.g., “rug”), we select the subset of neurons labeled with this concept, denoted as $\mathcal{N}_{l_{k}}\subset\mathcal{N}_{\mathcal{L}}$. These neurons are responsible for encoding the target concept.
@@ 4 Neuron-wise Representation Analysis
In this section, we conduct a neuronal analysis on infant models, focusing on the validation of beyond vocabulary neurons’ visual concepts using the proposed NeuronClassifier framework. This analysis aims to demonstrate the infant model’s strong generalization potential to unseen classes and, further, its capacity to capture higher-level visual concepts.
@@ 4.1 Setup
  • Class name in Konkle object dataset: All class labels from the Konkle object dataset are included.
@@
  • Infant-like models: The CVCL-ResNeXt50 model is trained on infant data, incorporating both egocentric frames and transcribed parent speech. To further test our hypothesis on visual concept acquisition in infant models, we also use DINO-S-ResNeXt50 [30], trained with the DINO [6] self-supervised approach on same dataset as CVCL. However, unlike CVCL, DINO-S-ResNeXt50 is trained only using video frames to get visual embeddings, without transcribed parent speech.
  • Large-scale pre-trained models: To establish an upper bound, we include two well pre-trained models, CLIP [32] and ResNeXt [40]. Although CLIP uses a ResNet50-based vision encoder rather than ResNeXt, we select CLIP-ResNet50 due to the architectural similarity between ResNeXt [40] and ResNet [19].
  • Baseline (randomized) model: As a lower bound, we introduce a randomized version of the CVCL-ResNeXt50 model. In this setup, the convolution layer weights in the vision encoder are initialized using Kaiming Initialization [18], with batch normalization weights set to one and all biases initialized to zero.
@@ 4.2 Results
In this section, we present the results of our NeuronClassifier framework applied to different models, especially CVCL. Our results demonstrate that the proposed framework effectively uncovers meaningful neuron that go beyond the model’s training vocabulary. Moreover, these findings, aligns with cogntitive perspective of vocabulary acquisition. We analyze the models’ performance across different $n$-way settings. Our method results in strong out-of-vocabulary classification as well as improved in-vocabulary classification.
@@
Figure 5: Class label coverage in visual concept neurons from various models. Percentage of Konkle object dataset [22] class labels discovered by our neuron labeling. In-vocabulary and out-of-vocabulary classes are defined by the CVCL model’s training vocabulary. Pre-trained models like CLIP and ResNeXt cover over half of the labels, while infant-trained models such as CVCL and DINO-S-ResNeXt [6] cover slightly fewer. CVCL-Randomized serves as a lower bound.
How well do the class corresponding visual concept neurons identified through neuron labeling (e.g. ”rug” as class name, rather than general discription as ”red”)? Figure 5 shows the percentage of number of classes in the Konkle object dataset that are discovered in visual concept neurons from each model during neuron labeling process. The results indicate that, well pre-trained models, such as CLIP-ResNet50 and ResNeXt50, demonstrate broader class name coverage. While CVCL perform slightly weaker, but still maintain around slightly less than half 50%. CVCL-Randomized is only around 28%. This class coverage reflects the models’ capacity to directly class corresponding meaningful representations during the neuron labeling process.
Figure 6: Age of acquisition (AoA) ratings for in-vocabulary and out-of-vocabulary words. Box plots illustrate the distribution of AoA ratings, while scatter points represent individual visual concepts from neurons in the CVCL model. Out-of-vocabulary words are acquired later than in-vocabulary words, indicating that the infant model captures higher-level visual concepts beyond its parental training vocabulary.
Age of acquisition ratings.
@@
We used AoA ratings from a dataset compiled by Kuperman, Stadthagen-Gonzalez, and Brysbaert  [24], which includes norms for over 30,000 English words gathered via Amazon Mechanical Turk. Each participant estimated the age in years at which they believed they first understood each word, even if they didn’t actively use them. This dataset is comparable to previously reported AoA norms gathered in laboratory settings.
Using this set of AoA norms, we compared mean AoA between in-vocab and out-of-vocab words discovered in CVCL’s internal representations. As shown in Figure 6, we found a significant difference between in-vocab and out-of-vocab AoA rating ( t(82) = 4.64, p $<$ 0.0001), in-vocab words (mean AoA = 4.99) are learned earlier than out-of-vocab words (mean AoA = 6.82). This pattern suggests that: (1) both sets of words are learned quite early, around later preschool and school years, with or without supervised labeling; (2) the difference in AoA between in-vocabulary and out-of-vocabulary words indicates that the infant model has developed a basic understanding of higher-level concepts (reflected by larger AoA). This foundational knowledge may lay the groundwork for word learning once corresponding parental speech is introduced.
@@
Table 1: Performance comparison in $4$-way classification. We leveraged visual concept neurons hidden in the representations of various models 4.1 to perform the same $4$-way classification in the previous work [38]. We discovered CVCL’s broader recognition ability, resulting in strong performance on out-of-vocabulary classes and significantly improving upon the vanilla approach in Figure 2. Here, “in-vocab” and “out-of-vocab” classes are defined globally based on the CVCL model’s vocabulary, with “all” representing the combined performance on both types.
@@
We evaluated the hidden potential of infant models’ vision encoders by applying our NeuronClassifier framework. Despite being trained on limited infant egocentric data, CVCL-ResNeXt50 demonstrates the ability to recognize similarly as nature of infant learning, achieving strong out-of-vocabulary classification performance. This result suggests that this infant-like model developed broader visual concepts that extend beyond linguistic input, similar to how infants naturally learn.
Additionally, we applied our method to in-vocabulary classification, where it outperformed the vanilla method previously used in CVCL [38], as introduced in Figure  2. Here, “in-vocab” and “out-of-vocab” classes are defined based on the CVCL model’s vocabulary, with “all” representing the combined performance on both types.
Notably, the DINO-S-ResNeXt50 [30], another infant model trained on the same dataset but without linguistic input, achieved similar performance using our method. This further proves that infant-like models have broader visual representations from egocentric visual inputs. Models such as ResNeXt50 and DINO-S-ResNeXt50 lack a text encoder and therefore cannot perform out-of-domain dataset classification without additional training; ours relies purely on their visual representations to perform classification.
Interestingly, under the vanilla evaluation on the Konkle object dataset, DINO-S-ResNeXt50 required additional fine-tuning for this downstream task. However, when using our method, it performed comparably to CVCL, highlighting the effectiveness of our approach. On the other hand, applying our NeuronClassifier framework to the original CLIP’s ResNet50 resulted in lower accuracy than using it directly. While the aim of this paper is not to improve the accuracy of neuron labeling methods, our goal is to discover the visual concepts learned by infants that extend beyond linguistic training input.

### v4

@@ Abstract
Infants develop complex visual understanding rapidly, even preceding of the acquisition of linguistic skills. As computer vision seeks to replicate the human vision system, understanding infant visual development may offer valuable insights. In this paper, we present an interdisciplinary study exploring this question: can a computational model that imitates the infant learning process develop broader visual concepts that extend beyond the vocabulary it has heard, similar to how infants naturally learn? To investigate this, we analyze a recently published model in Science by Vong et al., which is trained on longitudinal, egocentric images of a single child paired with transcribed parental speech. We perform neuron labeling to identify visual concept neurons hidden in the model’s internal representations. We then demonstrate that these neurons can recognize objects beyond the model’s original vocabulary. Furthermore, we compare the differences in representation between infant models and those in modern computer vision models, such as CLIP and ImageNet pre-trained model. Ultimately, our work bridges cognitive science and computer vision by analyzing the internal representations of a computational model trained on an infant visual and linguistic inputs. Our code is available at https://github.com/Kexueyi/discover_infant_vis.
@@
Figure 1: Inspired by the development of the infant visual system, CVCL [43] is trained on infant egocentric frames and transcribed parental speech (a) and demonstrates object recognition ability within the vocabulary provided by parental speech (c). However, infant visual development is not limited to parental guidance. Thus, we hypothesize that a computational model trained on an infant’s daily experiences can similarly acquire visual concepts beyond its training parental speech. To explore this, we perform neuron labeling (b) to identify visual concept neurons, including concepts that were never mentioned in the parental speech vocabulary (e.g., “rug”).. Based on discovered neurons, we show that the model can recognize objects beyond the training vocabulary, akin to early infant visual development (d). Some images are cited from prior work [43] as the original data is not accessible to us.
@@ 1 Introduction
Infants are remarkable learners, sparking interest across various academic disciplines. Computer vision is no exception, with researchers studying infant visual learning from various perspectives [1, 34, 37, 33, 43]. A recent milestone is Child’s View for Contrastive Learning (CVCL) by Vong et al.[43], which trains a model from scratch on longitudinal egocentric videos of a single infant (6–25 months) paired with transcribed parental speech to learn visual-linguistic associations, similar to CLIP [36]. The resulting model develops object recognition abilities, aligning with developmental psychology findings that infants acquire object names by linking words to visual referents [3, 18]. In this paper, we analyze the internal visual representations of CVCL to better understand its mechanism, similar to how developmental studies observe infants’ internal neurons [40].
In developmental psychology, research on real infants suggests that their object recognition is deeply influenced by their visual experiences with the world. A headcam study of 8.5 to 10.5-month-old infants [8] showed that the first nouns infants acquire often correspond to objects they see most frequently. While infants may also hear the names of these frequently seen objects, their early visual familiarity plays a critical role in object recognition. Visual understanding may develop before the learning of corresponding names, potentially facilitating the process of word acquisition [35]. For instance, newborn infants have shown an innate ability to recognize visual patterns [14], and their development of visual concepts often precedes the emergence of verbal thought [29].
Our Hypothesis: Based on these infant studies, we hypothesize that a computational model trained on an infant’s daily experiences may similarly acquire visual concepts extending beyond its linguistic training data. While CVCL demonstrates visual-linguistic mappings of objects named in parental speech, its visual recognition capability should not be limited to parental vocabulary supervision. We conjecture that the vision encoder may develop the ability to recognize concepts beyond these linguistic training data, similar to how infants form object recognition before learning object names.
@@
To investigate this, we analyze CVCL’s internal representations using network dissection [2, 32], or more intuitively, neuron labeling. Furthermore, we implement a neuron-based, training-free classification framework (NeuronClassifier) that leverages visual concepts identified via neuron labeling. Relying solely on these neurons, this approach not only achieves better recognition performance than originally reported [43], but also discovers internal visual concepts extending beyond the model’s training vocabulary, supporting our aligned hypothesis.
From a developmental psychology perspective, the visual concepts discovered beyond the training vocabulary tend to have a higher age of acquisition (AoA) [28] values (Figure 7). This results supports CVCL’s ability to develop visual understanding that precedes explicit labeling, mirroring cognitive studies where infants develop pre-verbal visual concepts [29]. This likely reflects real-world learning dynamics: children first acquire labels for frequent, concrete concepts (captured in CVCL’s training vocabulary), while visually grounded representations for unlabeled concepts still form earlier than their eventual linguistic acquisition.
@@
To further explore CVCL’s internal representations in the context of computer vision, we compare its visual features with widely-used representations such as CLIP and ImageNet. While CVCL is trained on a unique infant dataset with limited exposure to diverse scenes and a much smaller dataset compared to CLIP, it exhibits similar low-level features in its early layers. However, its higher-level features in the final layer differ significantly. These differences also extend to the visual concept neurons in the model’s deeper layers.
@@
  • We show that infant model have developed understanding beyond linguistic training inputs, by discovering visual concepts hidden in the model representations, aligning with existing infant studies.
  • We demonstrate that the discovered visual concept neurons can improve object recognition performance by implementing a training-free framework (NeuronClassifier).
  • We find that the infant model shares similar low-level representations with ImageNet or CLIP models but diverges in deeper layers due to a lack of diverse higher-level visual concepts.
@@ 2 Related Work
Modeling how children learn has long been a strategy for advancing artificial intelligence. Instead of directly replicating adult intelligence, Alan Turing suggested, “why not rather try to produce one which simulates the child’s?” [42] Training models on egocentric videos or multimodal data captured from infant perspectives aligns with this idea. [1, 33, 37, 34, 41, 43], as these videos approximate the input available to human infants during development. Our study established on CVCL [43], which we explain in Section 2.1.
@@
Our goal is to understand the model internal neurons trained on infant data. In this context, techniques for interpreting intermediate representations in deep neural networks are relevant. Beginning with Network Dissection [2], a method that quantifies alignment between hidden neurons and visual concepts, numerous studies have aimed to make black-box models more transparent. These methods enable compositional concept discovery [30], assignment of compositional concepts with statistical quantification [5], open-vocabulary neuron captioning [23], and the use of CLIP’s rich embeddings for neuron-concept alignment [32, 24]. In addition to direct neuron dissection, other approaches analyze component functions by decomposing image representations to reveal the role of attention heads within multimodal embedding spaces [16] or identifying neurons with similar functions across a diverse model zoo [12].
2.1 Preliminary: CVCL
@@
CVCL is trained on the SAYCam-S dataset [39], a longitudinal collection of egocentric recordings from a child aged 6 to 25 months, containing around 200 hours of video. To create meaningful image-text pairs for model training, transcripts were pre-processed to retain only child-directed utterances, excluding the child’s own vocalizations. Frames were extracted to align with utterance timestamps. The resulting dataset comprises 600,285 frames paired with 37,500 transcribed utterances, forming a multimodal dataset that simulates the sparse and noisy real-world experiences from which children learn.
@@
Employing a self-supervised contrastive learning approach akin to CLIP [36], CVCL learns to align egocentric visual frames with transcribed parental speech. Co-occurring pairs are treated as positive examples, while non-co-occurring pairs serve as negatives. This method allows the model to develop multimodal representations without external labels, imitating a child’s natural learning process.
@@
For evaluation, CVCL adopts a $n$-way classification task (Figure 2) in which the model selects the most relevant visual reference from a set comprising one target image and $n-1$ foil images. This approach is inspired by the intermodal preferential looking paradigm (IPLP) [17, 18], used in infant recognition studies to measure language comprehension through differential visual fixation. By aligning its visual and text encoders, CVCL achieves comparable in-domain test accuracy to models like CLIP. However, CVCL demonstrates relatively weak test performance on the Konkle object dataset [25], which includes naturalistic object categories on a white background, using only classes available in the training data.
@@ 3 Method
In this section, we describe how to explore neuron-level concepts and leveraging them in $n$ way classification tasks. We begin by using published neuron labeling techniques to discover visual concepts hidden within CVCL [43]. Then, we introduce our framework to utilize these labeled neuron concepts for $n$-way classification.
@@ 3.1 Neuron Labeling
We follow CLIP-Dissect [32] for internal representation analysis due to its flexibility in concept sets and input image dataset.
@@
$l_{k}=\arg\max_{m}\texttt{sim}(t_{m},q_{k};P),$ (3)
where $\texttt{sim}(\cdot,\cdot)$ represents the similarity function (e.g. cosine similarity or Soft-WPMI [44, 32]), and $t_{m}$ denotes the most similar text concept. Collecting the label $l_{k}$ for each neuron, we define the label vector $\mathcal{L}=[l_{1},l_{2},\dots,l_{K}]$ to represent the assigned concepts across the entire model.
During this neuron labeling process, we aim to assign meaningful concepts to each neuron. We use CLIP-dissect because:
  • Concept set $\mathcal{S}$: Instead of allowing an infinite range of possible concepts for neuron labeling, using a fixed concept set narrows this process by constraining it to a limited selection of concepts. Including diverse concepts in this set enables us to identify neurons corresponding to these visual concepts. This setup ensures that each neuron is assigned a specific label, though some labels may be spurious, as illustrated in Figure 3.
@@
Figure 3: Spurious labeled neurons in infant model CVCL by CLIP-Dissect, showing top-3 activated Konkle dataset images. This indicates: (1) not all neurons interpretable with human semantic concepts [13]; (2) neuron-labeling can produce spurious labels. To further support our hypothesis, we implement NeuronClassifier.
@@
The flexibility of the concept set $\mathcal{S}$ allowing us to discover visual concepts that the infant model has never encountered in its training data. In this process, CLIP serves as a well-pretrained miner, utilizing its rich image-text embeddings to identify concepts hidden within CVCL. By aligning the activations of CVCL’s neurons with CLIP’s embeddings, we discover meaningful hidden visual concepts within the infant model, even extending beyond the model’s linguistic training data. This approach leverages the diverse vocabulary of the concept set and the rich embeddings of CLIP to reveal visual concepts embedded in CVCL’s internal representations, each associated with corresponding neurons.
@@ 3.2 Neuron-Based Classification
How do we ensure that neurons representing concepts beyond the model’s original vocabulary truly exist within the network? In this section, we propose NeuronClassifier, a training-free framework that leverages neuron activations to detect and validate such concepts. By discovering neurons with specific visual concept, we aim to confirm the presence of these latent, beyond-vocabulary neurons and use them to perform $n$-way classification. The framework, illustrated in Figure 4, involves three main steps.
@@
Given a target label $l_{k}\in\mathcal{L}$, where $\mathcal{L}$ represents all neuron labels in the model, same as label vector in Section 3.1 (e.g., “rug”), we select the subset of neurons labeled with this concept, denoted as $\mathcal{N}_{l_{k}}\subset\mathcal{N}_{\mathcal{L}}$. These neurons are responsible for encoding the target concept.
@@ 4 Neuron-wise Representation Analysis
In this section, we conduct a neuron-wise analysis of infant models and provide implementation details. First, we perform neuron labeling on the infant CVCL model. Then, we use our NeuronClassifier framework to leverage visual concept neurons identified through neuron labeling, resulting in better recognition performance than the original approach [43] while also discovering internal visual concepts beyond the model’s training vocabulary. These results support our hypothesis.
@@ 4.1 Setup
  • Class in Konkle object dataset: All class labels from the Konkle object dataset are included.
@@
  • Infant models: CVCL is trained on unique infant data, using both egocentric frames and transcribed parent speech. We also take DINO-S-ResNeXt50 [33] as reference model compare recognition ability derived from same visual experience. It trained with DINO[6] self-supervised approach on the infant same dataset.
  • Broadly-trained models: To establish an upper bound, we include CLIP [36] and ResNeXt50 [45], which are broadly trained on large scale Internet images. Although CLIP uses a ResNet50-based vision encoder rather than ResNeXt, we select CLIP-ResNet50 due to the architectural similarity between ResNeXt [45] and ResNet [22].
  • Randomized model: As a lower bound, we introduce a randomized version of the CVCL-ResNeXt50 model. In this setup, the convolution layer weights in the vision encoder are initialized using Kaiming Initialization [21].
@@ 4.2 Results
In this section, we present our results and findings from applying neuron labeling and our NeuronClassifier framework to the infant CVCL model, compared with other reference models. We define two class types for our analysis:
  • In-vocabulary classes: Object classes present in the model’s training linguistic input, which also appear in test object class and are detectable in internal representations.
  • Out-of-vocabulary classes: Object classes not included in the training linguistic input but detected in test object class and internal representations.
Figure 5: In-vocabulary and out-of-vocabulary relationships visualized using a Venn diagram.
Our results demonstrate that the proposed framework effectively discovers meaningful neurons that represent concepts beyond the model’s training vocabulary. These findings align with the cognitive perspective of vocabulary acquisition in infant development. We analyze the models’ performance across different $n$-way classification settings, showing that our method yields strong out-of-vocabulary classification performance while simultaneously improving in-vocabulary classification accuracy.
@@
How well do visual concept neurons identified through neuron labeling correspond to specific class names (e.g., “rug”) rather than general descriptive attributes (e.g., “red”)? Figure 6 shows the percentage of classes in the Konkle object dataset that are discovered in visual concept neurons from each model during the neuron labeling process. The results indicate that well-pretrained models, such as CLIP-ResNet50 and ResNeXt50, demonstrate broader class name coverage. While CVCL performs slightly weaker, it still maintains coverage of slightly less than 50%. In contrast, CVCL-Randomized achieves only around 28% coverage. This class coverage metric reflects the models’ capacity to form class-corresponding meaningful representations during the neuron labeling process.
Figure 6: Class coverage in visual concept neurons. Percentage of Konkle object dataset [25] classes identified through neuron labeling. Broadly pre-trained models (CLIP, ResNeXt) achieve over 50% coverage, while developmentally inspired infant models (CVCL [43], DINO [6]) show lower coverage. CVCL-Randomized provides a lower-bound comparison.
Age of Acquisition (AoA) ratings.
@@
We used AoA ratings from a dataset compiled by Kuperman, Stadthagen-Gonzalez, and Brysbaert  [27], which includes norms for over 30,000 English words gathered via Amazon Mechanical Turk. Each participant estimated the age in years at which they believed they first understood each word, even if they did not actively use them. This dataset is comparable to previously reported AoA norms [38] gathered in laboratory settings.
Figure 7: Age of acquisition (AoA) ratings for in- and out-of-vocabulary visual concepts. Comparison of word acquisition timing [27] between in-vocabulary and out-of-vocabulary concepts. Out-of-vocabulary concepts tend to have a higher estimated acquisition ages in the infant-inspired CVCL model, indicating development of visual understanding beyond explicit linguist inputs.
Using this set of AoA norms, we compared mean AoA between in-vocab and out-of-vocab words discovered in CVCL’s internal representations. As shown in Figure 7, we found a significant difference between in-vocab and out-of-vocab AoA rating ( t(82) = 4.64, p $<$ 0.0001), in-vocab words (mean AoA = 4.99) are learned earlier than out-of-vocab words (mean AoA = 6.82). This pattern suggests that: (1) both sets of words are learned quite early, around later preschool and school years, with or without supervised labeling; (2) the difference in AoA between in-vocabulary and out-of-vocabulary words indicates that the infant model has developed a basic visual understanding of concepts with higher AoA. This foundational knowledge may lay the groundwork for word learning once corresponding parental speech is introduced.
@@
We evaluated the hidden potential of infant models’ vision encoders by applying our NeuronClassifier framework, summarized in Table 1. Despite being trained on infant egocentric data with limited amount and diversity, CVCL demonstrates the ability to recognize similarly as nature of infant learning, revealing strong out-of-vocabulary classification performance. This result suggests that this infant model developed broader visual concepts that extend beyond linguistic input, similar to how infants naturally learn. We also applied our method to in-vocabulary classification, where it outperformed the vanilla method previously used in CVCL [43], as introduced in Figure  2. “All” representing the combined performance on both “in-vocab” and “out-of-vocab”. These results support our hypothesis.
We include additional model comparisons: (1) The DINO-S-ResNeXt50 infant model [33], trained on the same dataset without text supervision, achieves comparable performance in visual representations. This suggests that models trained on identical data distributions with different self-supervised methods may yield similar representational outcomes. (2) Broadly-trained models such as ResNeXt50 and CLIP establish performance upper bounds. However, CLIP underperforms in the vanilla setting, relying exclusively on visual neurons without text encoder guidance. While this work does not aim to advance zero-shot learning methods, it reveals visual concepts in infant models that emerge independently of linguistic inputs.
In classification on the Konkle object dataset, both DINO-S and ImageNet ResNeXt50 required fine-tuning for this task. Our framework enables neuron-based classification without downstream fine-tuning, and providing a training-free qualitative inspection of internal representations.
Table 1: Neuron-based classification results in $4$-way evaluation among models in Section 4.1. “Vanilla” refers to classification based on image-text similarity (Figure 2). “” denotes cases where direct classification on the Konkle dataset [25] is not possible due to missing text encoder or need fine-tuning. By leveraging neurons discovered in the representation, NeuronClassifier enables broader recognition, particularly in CVCL (bolded for emphasis), achieving improved recognition in both in-vocabulary and out-of-vocabulary, supporting our hypothesis. “All” represents the combined performance on both in- and out-of-vocabulary.